OpenStudy (anonymous):

Mia wants to deposit $3,886 into savings accounts at three different banks: Bank of the US, 7/4 Bank, and Catch Bank. She will deposit three times as much into Bank of the US as 7/4 Bank because they offer a higher interest rate. She also expects the Catch Bank deposit to be 45% of the total of the other two deposits. How much will Mia deposit into the Catch Bank savings account? $670 $1,206 $1,340 $2,010

OpenStudy (anonymous):

u = bank of the u.s s = 7/4 bank c = catch bank u + s + c = 3886 u = 3s c = .45(u + s) c = .45(3s + s) c = .45(4s) c = 1.8s now sub known values into 1st equation u + s + c = 3886 3s + s + 1.8s = 3886 5.8s = 3886 s = 3886/5.8 s = 670 u = 3s u = 3(670) u = 2010 u + s + c = 3886 2010 + 670 + c = 3886 2680 + c = 3886 c = 3886 - 2680 c = 1206 check... u = 3s 2010 = 3(670) 2010 = 2010 (correct) or... c = 1.8s 1206 = 1.8(670) 1206 = 1206 (correct) or... u + s + c = 3886 2010 + 670 + 1206 = 3886 3886 = 3886 (correct) Catch Bank deposit is $ 1206

OpenStudy (anonymous):

THANK YOU SOOO MUCH!!

OpenStudy (anonymous):

its ok.....I don't mind helping out :)

OpenStudy (anonymous):

Can you help me with this one please? Edward received a gift card for $85 to a movie theater. He may use the gift card for either admission tickets or snacks. Tickets cost $7.50 and his favorite snack costs $3.75. If he wants to use his gift card for at least 15 individual purchases, which is a combination of tickets and snacks Edward may purchase? 7 tickets and 8 snacks 10 tickets and 6 snacks 8 tickets and 6 snacks 11 tickets and 3 snacks

OpenStudy (anonymous):

t stands for tickets s stands for snacks t + s = 15 --> s = 15 - t 7.5t + 3.75s = 85 now sub 15 - t in for s in the 2nd equation 7.5t + 3.75s = 85 7.5t + 3.75(15 - t) = 85 7.5t + 56.25 - 3.75t = 85 3.75t = 85 - 56.25 3.75t = 28.75 t = 28.75/3.75 t = 7.6 rounded down to 7 sub in 8 for t in the 1st equation t + s = 15 7 + s = 15 s = 15 - 7 s = 8 7 tickets and 8 snacks